A wooden beam with a rectangular section 150mm wide and 250 deep can resist a 140 kN-m bending stress. Determine the maximum compressive and tensile stress in the section if the beam material is non-isotropic with the tensile modulus of elasticity twice the compressive modulus.  



1
Expert's answer
2021-06-29T00:23:52-0400

Ultimate bending moment = "1.5*140 = 210 kN.m"

"I= \frac{bd^3}{12}= \frac{150*250^3}{12}=19.53*10^7"

"y=\frac{d}{2}=\frac{250}{2}=125 mm"

"=\sigma_{max}=\frac{My}{I}=\frac{210*10^6*125}{19.53*10^7} = 134.41 N/mm^2"

Hence compressive stress = "-134.41 N\/mm^2"

Hence tensile stress = "134.41 N\/mm^2"